Research discovered that the average heart rate of a sweeper in curling (a Winter Olympic sport) is 191 beats per minute. AssThe answers are in red. Please break it down for me, Thanks in advance!
Answer #1

5 solution Here mean (M) = 191, S.D() = 5 a here zun-u - 194-191 = 0.6 for a=194, 2= 5. P(x7194)=P(2706) = 1-p/2206) = 1-0-72